A compressed cylinder of gas contains 1.5 into ten to the power 3 g of N2 gas at a pressure of 2x10 to the power 7 pa and a temperature of 17.1 degree celsius. What volume of gas has been released into the atmosphere if the final pressure in the cylinder is 1.8 into ten to the power 5 pa? Assume ideal behaviour and that the gas temperature is unchanged. 

Grade:12

1 Answers

Vikas TU
3 years ago
Volume of nitrogen gas : V1 = 1.5*10^3 grams  (1g = 1mL)
Pressure of nitrogen gas ; P1 - 2*10^7 pascal
Volume of nitrogen gas at pressure P2 = V2
Pressure of nitrogen gas at volume  V2= P2= 1.8* 10^5  pascals
According to Boyle's law: P1v1 = P2v2
On solving for v2 , v2 = 1.66*10^5 mL